OpenStudy (anonymous):

F to the 3rd power over f to the 9th power

OpenStudy (cggurumanjunath):



OpenStudy (anonymous):

Law of Exponents: b^x/b^y = b^(x-y) So x^12/x^4 = x(12-4) = x^8

OpenStudy (anonymous):

Would it be f to the 6th

OpenStudy (cggurumanjunath):